OpenStudy (anonymous):

4 divided by 412 (not in decimals so if there is remander put R and then whatever!)

OpenStudy (anonymous):

103

OpenStudy (anonymous):

unless you mean \[\frac{4}{412}\] in which case it is \[\frac{1}{103}\]

OpenStudy (anonymous):

4/412 =1/103

OpenStudy (anonymous):

i mean 4/412 =___

OpenStudy (anonymous):

1/103

OpenStudy (anonymous):

so 103?? thxx!! :D

OpenStudy (anonymous):

u r dividing 4 by 103 so the answer is 1/103 not 103